SPECTROSCOPIC EVALUATION OF THE STRUCTURE AND DYNAMICS OF THE HYDROGEN BONDS IN DIFFERENT CRYSTAL MODIFICATIONS OF POTASSIUM BIIODATE

\underline{A.I.BARABASH}, T.A.GAVRILKO, G.A.PUCHKOVSKAYA, I.V.SEKIRIN, Institute of Physics, National Academy of Sciences of Ukraine, Prosp. Nauki, 46, 252650 Kiev, Ukraine; J.BARAN AND H.RATAJCZAK, Institute of Low Temperature and Structure Research, Polish Academy of Sciences, 2 Okolna Street, 50-950 Wroclaw, Poland.

The KIO_{3}·HIO_{3} crystals belong to the family of acidic iodate compounds which is well-known for its ferroelectric phase transitions and has four crystal modifications. Here we report the results of IR and Raman study of vibrational spectra (4000−60 cm^{−1}) of α and δ-modifications in the temperature range from 300 to 13 K. The α-KIO_{3}·HIO_{3} crystal (P 2_{1}/b space group, z=4 at room temperature) is built of two nonequivalent sets of dimers IO_{3}⋯H⋯O_{3}I, joined by strong symmetric hydrogen bond (O⋯O distance is about 2.5 Å). At 223 K (T_{c}) the ferroelectric phase transition occurs in this crystal. The stable δ-modification (space group F dd2, z=24 at room temperature) is formed from structure units of three distinct types: I_{2}O_{5} and I_{2}O_{5}(OH), which are joined by the H-bond network. In order to understand the normal modes of vibrations in α-KIO_{3}·HIO_{3} and δ-KIO_{3}·HIO_{3}, we have carried out lattice dynamics calculations using a generalized valence force field for the dimers [H(IO_{3})_{2}]^{−} and other structure units. The introduction of the translation and libration coordinates related to the normal coordinate analysis leads to a precise assignment of the frequencies and shows that both stretching and bending vibrations of IO_{3} groups are strongly coupled with the translational vibrations of the hydrogen atom in case of αmodification. In the case of δ-modification those vibrations are localized on separate I−O bonds. A comparative analysis of the vibrational spectra and structure is presented. On the base of analysis of vibrational spectra and according to measurements^{a} of pressure dependencies of the ^{127}I NQR spectra at 77 K it was concluded that the mechanism of ferroelectric phase transition in α-modification of potassium biiodate crystal involves the proton ordering in the double-well potential at T_{c}.

aA.I.Barabash, J. Mol. Struct., \underline{270}, 517 (1992).